given a 3 percent interest rate, compute the year 6 future value of deposits made in years 1,2,3 and 4 of $1350 $1650 $1650 and 1750 respectively
Answer #1

Rate = 3%

If deposits are made at the beginning of years

FV of Deposit 1 = 1350*(1+3%)^6 = 1611.97

FV of Deposit 2 = 1650*(1+3%0^5 = 1912.80

FV of Deposit 3 = 1650*(1+3%)^4 = 1857.09

FV of Deposit 4 = 1750*(1.03)^3 = 1912.27

If deposits are made at the end of years

FV of Deposit 1 = 1350*(1+3%)^5 = 1565.02

FV of Deposit 2 = 1650*(1.03)^4 = 1857.09

FV of deposit 3 = 1650*(1.03)^3 = 1802.99

FV of Deposit 4 = 1750*(1.03)^2 = 1856.57